OpenStudy (anonymous):

slove for y and show your work: 3y+2x=16

OpenStudy (anonymous):

3y+2x-2x=16-2x 3y=16-2x /3 /3 /3 y=(16-2x)/3

OpenStudy (lalaly):

3y=16-2x y=(16-2x)/3